# Better Together? A Field Experiment on

### v2

@@ 4.1 Impact of Algorithm Availability on Child Health and Wellbeing
Importantly, tool access did not crowd out resources provided to children in the control group. Table A.7 tests for the presence of short-run, within-day spillovers (i.e., contamination) by leveraging the share of referrals that were treated during the day, excluding a child’s own referral. If benefits had come at the expense of the control group, then, assuming fixed resources per day, one might expect to see: (1) worse outcomes for children in the control group as the percent treated in the day increases, and possibly also (2) weaker effects for treated children as the percent of children treated in the day increases. We find no evidence of spillover effects in terms of screen-ins, discussion time, or child harm. Children for whom the tool was available benefited without imposing costs on other children.
@@
Our specification may underestimate the benefits of the algorithm if workers learned and changed their investigation tendencies from using the tool. In this scenario, we would expect indirect benefits for children in the control group and an attenuation of treatment effects over time. Splitting children into three equally-sized groups by referral date, we find no evidence that workers learned from the algorithmic tool over the course of the trial (Figure A.7). This is consistent with anecdotal evidence from interviews with workers. To the extent that there was unobserved learning by workers, our study estimates would likely be a conservative lower bound on the true impacts of providing the algorithm.
@@ 4.2 Impact on Group-Level Disparities in Health and Wellbeing
In this section, we provide empirical evidence of how algorithm support affects disparities. A significant concern about algorithmic decision aids is their potential to perpetuate and widen inequality. Algorithms trained on past decisions could embed historical bias (Obermeyer et al., 2019) or be less informative for members of disadvantaged groups (Larrazabal et al., 2020; Blattner and Nelson, 2022). On the other hand, algorithms may be less biased than status-quo human decision-makers (c.f., Hoffman et al. 2018, Kleinberg et al. 2018, Arnold et al. 2024). They may also reverse bias depending on selection into the training data (Rambachan and Roth, 2020), reveal hidden disparities unknown to practitioners (Pierson et al. 2021), or cause people to think about information more deliberatively in a manner that reduces subconscious bias (Agan et al. 2022). Human use of algorithms also has a theoretically ambiguous impact on disparities, underscoring the need for evaluation.
@@
Notes: These figures show the mean value of children’s harm index by demographic category and tool availability, derived from regression estimates. SNAP (food stamp) receipt is a proxy for lower socioeconomic status. White children in Panel (i) have confidence intervals that are too small to be visible. 90% confidence intervals are reported and randomization controls are included throughout.
We report effects of algorithm availability on disparities in health by race, gender, ethnicity, and socioeconomic status (Figure 1 and Table A.8). When the tool was unavailable, as shown in light blue on the left side of each panel of Figure 1, children’s harm index was elevated for Black children, Hispanic children, girls, and children from low-income families as measured by SNAP (food stamp) receipt. Table A.8 (row “Type”) formalizes this result: in the control group, differences between the specified demographic group and out-group are statistically significant for girls, Hispanic children, and low-income children, and large but imprecise for Black children. These disparities mirror known statistics showing children with these demographic characteristics tend to face heightened risk of child mortality (Currie and Schwandt, 2016) and/or maltreatment (Putnam-Hornstein et al., 2021).
Access to the algorithm jointly reduced harm for children in disadvantaged categories, namely Black, Hispanic, female, and low-income children (compare light blue triangles to dark blue triangles in Figure 1). The out-groups (square dots) also experienced improvements in outcomes, but they were smaller in magnitude. Precise estimates are reported in Table A.8, and imply a reduction in disparities in the harm index on the order of 50 to 80 percent (Table A.8, bottom row). Although none of the reductions in disparities is statistically significant on its own, the cumulative evidence for a reduction in disparities is strong: we can confidently reject that all of the disparities are unaffected ($p<0.001$, Table A.8).
@@ 4.3 Impact on Racial Disparities in CPS Screen-Ins
Providing algorithmic information may affect disparities in agency surveillance. CPS investigations can sometimes be intrusive and traumatic, and prior work has documented substantial racial disparities – in some cases unwarranted – in CPS contact across the U.S. (Kim et al. 2017, Edwards et al. 2021, Baron et al. 2023). In our sample, Black children are nearly twice as likely to be investigated as other children, even after conditioning on algorithm risk score (Table A.9). In contrast, we find no evidence of screen-in disparities by gender, ethnicity, or socioeconomic status even after controlling for algorithm-predicted risk (Table A.9). We therefore focus our analysis of surveillance disparities on race, specifically Black children relative to children who are not Black (primarily white).
Figure 2 shows that, in the control group, Black children at below-median predicted risk were 60 percentage points more likely to be screened in than other (primarily white) children of similar risk level. However, when the algorithm was available, below-median-risk Black children were screened in at much lower rates, comparable to rates for other children. Column 2 of Appendix Table A.10 shows that the algorithm tool altogether reduced screen-ins by 13 percentage points for Black children across the entire sample, though point estimates are imprecisely estimated. These findings suggest that access to the algorithm reduced unwarranted racial disproportionality in investigations among low-risk children. Although the trial county has a relatively small Black population, there were still $n=170$ Black children in the full sample when using CPS-derived race alone, with at least 23 children in each race-by-treatment-status cell. Regression estimates and robustness to including a permutation test and alternate ways of defining race are reported in Appendix Table A.10. Although the exact magnitude of screen-in disparity reductions – and corresponding power to detect significant effects – depends on the specification, all estimates suggest the algorithm reduced surveillance disparities conditional on risk and, taking the imprecise point estimates at face value, may have canceled out disparities entirely. Though we are statistically underpowered, we do not detect an offsetting increase in the harm index among low-scoring Black children (Appendix Figure 9i).
@@ 5.1 Improved Worker Allocation of Screen-Ins
The intervention did not increase the total number of investigations (“screen-ins”) or any other measurable level of services (Figure A.10). Workers did not increase overall levels of intervention, did not significantly change discussion length, and did not provide greater levels of one intervention in place of another, as is sometimes observed due to asymmetric valuation of false positives and false negatives (Almog et al. 2024, Fitzpatrick et al. 2022). In our setting, levels of intervention may have been relatively inelastic given legal, financial, and time constraints.
@@
Furthermore, workers did not increase their rate of investigation for children with high algorithm risk scores. Figure A.11 shows that the relationship between screen-in rates and risk scores is modest in the control group, and changes little with the introduction of the tool. A caller’s current concerns about a child are often more determinative of the decision to screen in than family history. The absence of an increase in screen-ins for high-scoring children is consistent with the idea that high-scoring children have often already been investigated (30 percent of children with the highest risk score, 20, were investigated in the year prior to their referral) and are not necessarily those who benefit the most from a new investigation. [^14] An earlier draft of this paper showed a small-in-magnitude, marginally significant ($p=0.098$) reduction in the screen-in by risk score gradient, when workers had access to the algorithm relative to the control group. This was due to the inclusion of $n=544$ children in our sample (those sent to a residual, overflow meeting group) where workers did not access the tool. Including this set of children introduced a misleading statistical anomaly, and our current sample excludes them. We have confirmed with text data from discussion notes that workers in these overflow groups did not consult the tool. Figure A.12 plots estimates from regressing other types of CPS interventions on algorithm access interacted with algorithm risk score. The tool did not increase the likelihood of other interventions for children with high algorithm scores. The only change is that workers appear to open more (fewer) prevention cases to children with low (high) algorithm scores, but these changes cannot explain the reduction in child harm.
@@
With algorithm support, however, workers did reallocate their investigations to children at higher risk of harm along other dimensions. Randomization allows us to infer improvements in decision-making by comparing realized outcomes for screened-out children in the intervention and control groups. Intuitively, levels of potential harm are equal in the algorithm and control groups prior to intervention, so a reduction in child harm for screened-out children implies a reallocation of screen-ins toward riskier children. This aligns with a CPS goal of determining which children are at low risk of harm and do not need to be visited. This analysis is our first and primary targeting test, reported in Column 1 of Table 3.
@@
The test in Column 1 is our best measure of targeting because it uses realized child harm, whether or not such harm could be predicted by the researchers. It captures the full measurable extent (given our main outcome) of the improvement in allocating investigations. When workers had access to the algorithm, the eventual realized harm index of children who were screened out on their initial referral was 0.074 standard deviations lower than when workers did not have the algorithm. Because CPS did not intervene for these children, differences in post-referral hospitalization outcomes reflect a difference in targeting alone. Workers therefore must have improved their assessment of which families were low risk and accordingly screened them out. Results are similar, with loss of precision, when restricting to children never screened in during the entire trial. This targeting improvement is even marginally more significant ($p=0.045$) after controlling for all the observables in our balance test, including harm prior to the referral and the number of prior hospital admissions. Algorithm support thus reallocated investigations away from lower-risk children, as measured by ex-post hospitalization. Figure 6, discussed later in Section 6.2, provides visual detail for this result.
In columns 2-4 of Table 3, we provide further evidence that access to an algorithm improved worker targeting of screen-ins to children in need. First, investigators were more likely to find injuries on screened-in children (Column 2). CPS-discovered injuries sustained prior to investigation increased by five percentage points (85 percent), a fifth of which were abrasions, black eyes, bone fractures, bruises, burns, cuts, scratches, and skull fractures. This variable was documented by CPS workers, rather than hospitals, providing an independent outcome margin measured for all screened-in children. The changes appear to reflect differences in the types of children screened in rather than downstream investigator behavior, as investigators were not informed of the algorithm score during their standard investigation protocol. Second, the algorithm tool helped workers target screen-ins to children who experienced more maltreatment-suggestive hospitalizations prior to the screen-in decision, despite teams not directly observing child medical history (Column 3). This result applies to the set of children in the highest quartile of predicted risk (scoring 16-20), which includes a significant majority of children with prior hospitalization. For children in the bottom three algorithm score quartiles, harm levels in the 180 days prior to the referral are too low to detect any effect. Finally, column 4 shows that workers increased screen-ins to children with a greater predicted harm index. Predicted harm is estimated using a rich set of characteristics described in the footnote of Table 3 and constructed using the leave-one-out sample procedure described in Abadie et al. (2018) to prevent overfitting predictions in the control group. Children screened in using the algorithmic tool had a tenth of a standard deviation (0.116 SD) greater predicted harm index than children screened in without algorithm support.
Table 3: Algorithm Tool Improved Targeting of Screen-Ins to Children
@@
Harm Index, Screened Out Children CPS Found Injury, Screened In Children Prior Harm Index, Top Quartile Algo Risk Score Predicted Harm Index
Algorithm Available -0.074* 0.051** -0.117 -0.008
@@
Algorithm Available $\cdot$ Screen-In 0.249* 0.116***
@@
Control Mean -0.001 0.060 0.146 0.006
Observations 2,368 1,065 1,320 3,412
Notes: This table reports four tests of whether the algorithm improved the targeting of investigations to children in the trial. Column 1 estimates the impact of tool availability on the realized harm index for children who CPS screened out during the child’s first referral in the trial. Column 2 estimates the effects of tool availability on the likelihood of CPS discovering a child injury, conditional on CPS screening the child in for investigation. Column 3 shows changes in targeting based on children’s harm index in the 180 days prior to their initial referral in the trial. This analysis is applied only to a subset of children comprising the top quartile of algorithm risk (score of 16-20). In Column 4, we regress a predicted harm index measure on an interacted model of screen-in and tool availability. Predicted harm is constructed using the leave-one-out procedure of Abadie et al. (2018) and predicts a child’s future harm index using features determined prior to the referral: family history prior to the referral, hospitalizations prior to the referral, demographics at the time of the referral, and basic referral characteristics. Estimates are restricted to the hospital outcomes sample, with a handful of observations omitted due to missing feature variable values. Randomization controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01.
@@
It may seem surprising that workers better allocated investigations based on potential child harm, despite not increasing investigations for children with high algorithm risk scores. These facts permit a number of hypotheses, and rule out others (Table A.11). The hypothesis most consistent with our findings is that workers leveraged expertise that was complementary to the algorithm, which we explore in greater detail in Section 5.3. Reminiscent of Bhatt et al. (2024), social workers in our setting may have allocated screen-ins partly based on expected treatment effects (proverbially $\hat{\beta}$) rather than expected outcome levels ($\hat{Y}(0)$). In other words, not all children with high algorithm scores will necessarily benefit from an additional CPS investigation, and vice versa.
Finally, the magnitudes for the improvement in screen-in targeting reported in columns 1 and 4 of Table 3 (approximately 0.1 SD) are large enough to potentially explain the entire reduction in child harm with algorithm support (-0.06 SD). Enhanced targeting of screen-ins could have also led to better targeting of subsequent CPS actions, though we are underpowered to detect this. Targeting of other concurrent interventions (e.g., law enforcement, community services) could also have improved, but screen-ins are the primary margin of CPS attention. In either event, child outcomes improved with algorithm support because workers made better assessments of child risk, as evidenced by better screening decisions. We provide additional evidence of this in Section 5.2.

### v3

@@ 4.1 Impact of Algorithm Availability on Child Health and Wellbeing
Importantly, we do not find evidence that tool access crowded out resources provided to children in the control group. Table A.8 tests for the presence of short-run, within-day spillovers (i.e., contamination) by leveraging the share of referrals that were treated during the day, excluding a child’s own referral. If benefits had come at the expense of the control group, then, assuming fixed resources per day, one might expect to see: (1) worse outcomes for children in the control group as the percent treated in the day increases, and possibly also (2) weaker effects for treated children as the percent of children treated in the day increases. Though our confidence intervals are large, we find no evidence of statistically significant spillover effects in terms of screen-ins, discussion time, or child harm. Overall, we interpret the results as suggestive, reassuring evidence that the algorithm benefited children in the treatment group without imposing costs on the control group.
@@
Our specification may underestimate the benefits of the algorithm if workers learned and changed their investigation tendencies from using the tool. In this scenario, we would expect indirect benefits for children in the control group and an attenuation of treatment effects over time. While our study is unable to assess learning that might have occurred after the end of the trial, we can observe differences in treatment effects for children referred at different points during the trial. Splitting the sample into four equally-sized groups by date of first referral in the trial, we find that harm is reduced by similar amounts for children referred in the first three quarters of the trial. There appears to be no impacts for children referred in the final quarter (Figure A.7), possibly owing to mechanical attenuation from a limited window to observe hospital outcomes for treated versus control during the trial (see footnote 13). The stability of estimates for the first three quarters of the trial are consistent with a lack of learning. However, these results should be interpreted with caution given the large confidence intervals by quarter and the fact that chronically-referred children were more likely to appear in earlier quarters of the trial. Although we cannot rule out learning within the trial, anecdotal evidence from interviews with workers suggest that workers did not consciously change their investigation strategy for control referrals. To the extent that there was unobserved learning by workers, our study estimates would likely be a conservative lower bound on the true impacts of providing the algorithm. Our discussions with supervisors at the end of the trial also suggest that workers did not become overly reliant on the tool, as supervisors correctly recalled that the tool only summarized history and not all dimensions of “future maltreatment risk.”
@@ 4.2 Impact on Group-Level Disparities in Health and Wellbeing
In this section, we provide empirical evidence of how algorithm support affects disparities. A significant concern about algorithmic decision aids is their potential to perpetuate and widen inequality. Algorithms trained on past decisions have been repeatedly shown to embed historical bias (Obermeyer et al., 2019) especially when historical data was from automatic decision-making (Agan et al., 2022). They may also be less informative for members of disadvantaged or under-represented groups (Larrazabal et al., 2020; Blattner and Nelson, 2022). Humans have been documented to override algorithm recommendations in ways that reinforce existing biases (Albright, 2019, e.g.). On the other hand, algorithms may be less biased than status-quo human decision-makers (c.f., Hoffman et al. 2018, Kleinberg et al. 2018, Arnold et al. 2024). Absence of rigorous empirical evidence on whether algorithmic tools indeed increase disparities has been a major hindrance to their adoption for social policy.
@@
Notes: These figures show the mean value of children’s harm index by demographic category and tool availability, derived from regression estimates. SNAP (food stamp) receipt is a proxy for lower socioeconomic status. White children in Panel (i) have confidence intervals that are too small to be visible. 90% confidence intervals are reported and randomization controls are included throughout. Regression estimates with p-values are provided in Table A.9.
We provide some of the first field experiment evidence of the impact of algorithm tools on disparities in a social policy setting. Specifically, we report effects of algorithm availability on disparities in health by race, gender, ethnicity, and socioeconomic status (Figure 1 and Table A.9). When the tool was unavailable, as shown in light blue on the left side of each panel of Figure 1, children’s harm index was elevated for Black children, Hispanic children, girls, and children from low-income families as measured by SNAP (food stamp) receipt. Table A.9 (row “Type”) formalizes this result: in the control group, differences between the specified demographic group and out-group are statistically significant for girls, Hispanic children, and low-income children, and large but imprecise for Black children. These disparities mirror known statistics showing children with these demographic characteristics tend to face heightened risk of child mortality (Currie and Schwandt, 2016) and/or maltreatment (Putnam-Hornstein et al., 2021).
Point estimates suggest that access to the algorithm reduced harm for children in disadvantaged categories, namely Black, Hispanic, female, and low-income children (compare light blue triangles to dark blue triangles in Figure 1). The out-groups (square dots) also experienced improvements in outcomes, but they were smaller in magnitude. Precise estimates are reported in Table A.9, and imply a reduction in disparities in the harm index on the order of 50 to 80 percent (Table A.9, bottom row). Though large, these reductions are not statistically distinguishable from zero. [^15] Earlier versions of the paper used Stata’s suest command, which reported a highly significant joint test $p<0.01$. However, suest does not accommodate different variables temporarily sharing the same name across specifications (i.e., a loop over four disadvantaged groups of interest), leading to an error when compared with hand-calculated stacked regression specifications.
Our findings are nevertheless informative in a Bayesian perspective. In the absence of prior evidence, suppose a policymaker would have assumed effects on disparities could be of any value. We model this with a uniform prior distribution where algorithm support could reduce disparities by up to 200 percent or increase them by that same amount. Following Brännlund et al. (2024), we then calculate the posterior distribution given the data we observe, assuming normally distributed point estimates with observed variance. [^16] The posterior is a truncated normal distribution over [-200,200]: $p(\theta|\hat{\theta})=\frac{1}{\Phi(\frac{200-\hat{\theta}}{\sigma})-\Phi(\frac{-200-\hat{\theta}}{\sigma})}\frac{1}{\sqrt{2\pi\sigma^{2}}}\exp\left(-\frac{(\hat{\theta}-\theta)^{2}}{2\sigma^{2}}\right)$ if $\theta\in[-200,200]$ and zero otherwise. $\Phi$ denotes a normal cumulative distribution function. We report these posterior distributions with 90 percent credible intervals in Appendix Figure A.8. Our estimates reveal that in most cases we can rule out a worsening of disparities for ethnicity, gender, and socioeconomic status.
@@ 4.3 Impact on Racial Disparities in CPS Screen-Ins
Providing algorithmic information may affect disparities in agency surveillance. CPS investigations can sometimes be intrusive and traumatic, and prior work has documented substantial racial disparities – in some cases unwarranted – in CPS contact across the U.S. (Kim et al. 2017, Edwards et al. 2021, Baron et al. 2023). Though a small share of the sample, Black children are nearly twice as likely to be investigated as other children in the sample, even after conditioning on algorithm risk score (Table A.10). In contrast, we find no evidence of screen-in disparities by gender, ethnicity, or socioeconomic status even after controlling for algorithm-predicted risk (Table A.10). We therefore focus our analysis of surveillance disparities on race, specifically Black children relative to children who are not Black (primarily white).
Figure 2 shows that, in the control group, Black children at below-median predicted risk were 60 percentage points more likely to be screened in than other (primarily white) children of similar risk level. However, when the algorithm was available, below-median-risk Black children were screened in at much lower rates, comparable to rates for other children. Column 2 of Appendix Table A.11 shows that the algorithm tool altogether reduced screen-ins by 13 percentage points for Black children across the entire sample, though point estimates are imprecisely estimated. These findings suggest that access to the algorithm reduced unwarranted racial disproportionality in investigations among low-risk children. Although the trial county has a relatively small Black population, there were still $n=170$ Black children in the full sample when using CPS-derived race alone, with at least 23 children in each race-by-treatment-status by above or below median risk cell. Regression estimates and robustness to including a permutation test and alternate ways of defining race are reported in Appendix Table A.11. Although the exact magnitude of screen-in disparity reductions – and corresponding power to detect significant effects – depends on the specification, all estimates suggest the algorithm reduced surveillance disparities conditional on risk and, taking the imprecise point estimates at face value, may have canceled out disparities entirely. Though we are statistically underpowered, we do not detect an offsetting increase in the harm index among low-scoring Black children (Appendix Figure 10i).
@@ 5.1 Improved Worker Allocation of Screen-Ins
The intervention did not increase the total number of investigations (“screen-ins”) or any other measurable level of services (Figure A.11). Workers did not increase overall levels of interventions and did not significantly change discussion length, as is sometimes observed due to asymmetric valuation of false positives and false negatives (Almog et al. 2024, Fitzpatrick et al. 2022). In our setting, levels of intervention may have been relatively inelastic given legal, financial, and time constraints.
@@
Notably, workers also did not increase their rate of investigation for children with high algorithm risk scores. Figure A.12 shows that the relationship between screen-in rates and risk scores is essentially flat in the control group and changes little with the introduction of the tool. [^17] An earlier draft of this paper showed a small-in-magnitude, marginally significant ($p=0.098$) reduction in the screen-in by risk score gradient, when workers had access to the algorithm relative to the control group. This was due to the inclusion of $n=544$ children in our sample (those sent to a residual, overflow meeting group) where workers did not access the tool. Including this set of children introduced a misleading statistical anomaly, and our current sample excludes them. We have confirmed with text data from discussion notes that workers in these overflow groups did not consult the tool. Although this may appear counterintuitive – especially because the algorithm risk score is strongly predictive of downstream CPS involvement and child harm – there are several reasons why a flat gradient can be consistent with both (i) workers making informed, non-random screening decisions and (ii) the tool providing meaningful predictive information. First, screening decisions are strongly influenced by the caller’s present concerns, including potential legal requirements for following up on certain concerns, which are often more determinative of an investigation than family history in isolation. Second, many high-scoring children already have ongoing cases and many have already recently been investigated (e.g., 30 percent of children with the highest algorithm risk score were investigated in the year prior to their referral), potentially reducing the marginal value of a new investigation for them. As a result, it is possible for the algorithm score to be informative about a child’s history and improve the allocation of investigations without necessarily increasing the slope. (In Section 6, we consider a counterfactual scenario where investigations are explicitly allocated to high-scoring children.) Finally, Figure A.13 plots estimates from regressing other types of CPS interventions on algorithm access interacted with algorithm risk score. The tool did not increase the likelihood of other interventions for children with high algorithm scores. The only change is that workers appear to open more (fewer) prevention cases to children with low (high) algorithm scores, but these changes cannot explain the reduction in child harm.
@@
With algorithm support, however, workers did reallocate their investigations to children at higher risk of harm along other (non-algorithm-score) dimensions. Randomization allows us to infer improvements in decision-making by comparing realized outcomes for screened-out children in the intervention and control groups. Intuitively, levels of potential harm are equal in the algorithm and control groups prior to intervention, so a reduction in child harm for screened-out children implies a reallocation of screen-ins toward riskier children. This aligns with a CPS goal of determining which children are at low risk of harm and do not need to be visited. This analysis is our first and primary targeting test, reported in Column 1 of Table 2.
@@
The test in Column 1 is our best measure of targeting because it uses realized child harm, whether or not such harm could be predicted by the researchers. It captures the full measurable extent (given our main outcome) of the improvement in allocating investigations. When workers had access to the algorithm, the eventual realized harm index of children who were screened out on their initial referral was 0.06 standard deviations lower than when workers did not have the algorithm. Because CPS did not intervene for these children, differences in post-referral hospitalization outcomes reflect a difference in targeting alone. [^18] Investigations are the primary institutional and measurable margin of intervention, though targeting of lower-level services could also have improved. Workers therefore must have improved their assessment of which families were low risk and accordingly screened them out. Results are similar, with loss of precision, when restricting to children never screened in during the entire trial. This targeting improvement is robust and marginally more significant after controlling for all the observables in our balance test, including harm prior to the referral and the number of prior hospital admissions. Algorithm support thus reallocated investigations away from lower-risk children, as measured by ex-post hospitalization. Figure 6, discussed later in Section 6.2, provides visual detail for this result.
Table 2: Algorithm Tool Improved Allocation of Screen-Ins to Children
@@
Harm Index, Screened-Out Children CPS Found Injury, Screened-In Children Prior Harm Index, Top Quartile Algo Risk Score Predicted Harm Index
Algorithm Available -0.060* 0.051** -0.123 -0.009
@@
Algorithm Available $\cdot$ Screen-In 0.258** 0.068*
@@
Control Mean 0.001 0.060 0.146 0.001
Observations 3,266 1,065 1,320 4,662
Notes: This table reports four tests of whether the algorithm improved the targeting of investigations to children in the trial. Column 1 estimates the impact of tool availability on the realized harm index for children who CPS screened out during the child’s first referral in the trial. Column 2 estimates the effects of tool availability on the likelihood of CPS discovering a child injury, conditional on CPS screening the child in for investigation. Column 3 shows changes in targeting based on children’s harm index in the 180 days prior to their initial referral in the trial. This analysis is applied only to a subset of children comprising the top quartile of algorithm risk (score of 16-20). In Column 4, we regress a predicted harm index measure on an interacted model of screen-in and tool availability. Predicted harm is constructed using the leave-one-out procedure of Abadie et al. (2018) and predicts a child’s future harm index using features determined prior to the referral: family history prior to the referral, hospitalizations prior to the referral, demographics at the time of the referral, and basic referral characteristics. Randomization controls are included throughout. Standard errors are clustered at the household level. Significance reported as: * p<<0.1, ** p<<0.05, *** p<<0.01. We report in Figures A.11 and A.12 that overall screen-in levels did not change.
In columns 2-4 of Table 2, we provide further evidence that access to an algorithm improved worker targeting of screen-ins to children in need. First, investigators were more likely to find injuries on screened-in children (Column 2). CPS-discovered injuries sustained prior to investigation increased by five percentage points (85 percent), a fifth of which were abrasions, black eyes, bone fractures, bruises, burns, cuts, scratches, and skull fractures. This variable was documented by CPS workers, rather than hospitals, providing an independent outcome margin measured for all screened-in children. The changes appear to reflect differences in the types of children screened in rather than downstream investigator behavior, as investigators were not informed of the algorithm score during their standard investigation protocol. Second, the algorithm tool helped workers target screen-ins to children who experienced more maltreatment-suggestive hospitalizations prior to the screen-in decision, despite teams not directly observing child medical history (Column 3). This result applies to the set of children in the highest quartile of predicted risk (scoring 16-20), which includes a significant majority of children with prior hospitalization. For children in the bottom three algorithm score quartiles, harm levels in the 180 days prior to the referral are too low to detect any effect. Finally, column 4 shows that workers increased screen-ins to children with a greater predicted harm index. Predicted harm is estimated using a rich set of characteristics described in the footnote of Table 2 and constructed using the leave-one-out sample procedure described in Abadie et al. (2018) to prevent overfitting predictions in the control group. Children screened in using the algorithmic tool had a tenth of a standard deviation (0.07 SD) greater predicted harm index than children screened in without algorithm support.
@@
It may seem surprising that workers better allocated investigations based on potential child harm, despite not increasing investigations for children with high algorithm risk scores. These facts permit a number of hypotheses, and rule out others (Table A.12). The hypothesis most consistent with our findings is that workers leveraged expertise that was complementary to the algorithm, which we explore in greater detail in Section 5.3. Reminiscent of Bhatt et al. (2024), social workers in our setting may have allocated screen-ins partly based on expected treatment effects (proverbially $\hat{\beta}$) rather than expected outcome levels ($\hat{Y}(0)$). In other words, not all children with high algorithm scores necessarily benefit from an additional CPS investigation, and vice versa.
Finally, the magnitudes for the improvement in screen-in targeting reported in columns 1 and 4 of Table 2 (0.06 SD) are large enough to potentially explain the entire reduction in child harm with algorithm support (-0.05 SD). Enhanced targeting of screen-ins could have also led to better targeting of subsequent CPS actions, though we are underpowered to detect this. Targeting of other concurrent interventions (e.g., law enforcement, community services) could also have improved, but screen-ins are the primary margin of CPS attention. In either event, child outcomes improved with algorithm support because workers made better assessments of child risk, as evidenced by better screening decisions. We provide additional evidence of this in Section 5.2.